OpenStudy (anonymous):

Subtract a-4 from 4a+16 and simplify your answer. The variable a should appear only once in your answer.

OpenStudy (anonymous):

4a+16-(a-4) =4a+16-a+4 =3a+20

OpenStudy (anonymous):

(4a +16) - (a-4)= 4a+16-a+4 = 3a+20

OpenStudy (uri):